OpenStudy (anonymous):

Given that F(x) = x² + 2, evaluate F(1) + F(5).

OpenStudy (ajprincess):

to find f(1) substitute x=1 in f(x). that is f(1)=(1)^2+2 to find f(5) substitute x=5 in f(x) that is f(5)=(5)^2+2 Nw add them together to get f(1)+f(5)

OpenStudy (anonymous):

F(x) = x² + 2 F(1) = (1)² + 2 = 1 + 2 = 3 F(5) = (5)² + 2 = 25 + 2 = 27

OpenStudy (anonymous):

Hence F(1) + F(5) = 3 + 27 = 30